If one of the roots of x2-bx+c=0 , (b,c belong to rational numbers) is (7-4(3)1/2)1/2 then :
(A) logbc=0
(B) b+c=5
(C) logcb=0
(D) bc=-4

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
when simplified
 
[7 – 4 sqrt(3)]^1/2 = 2 – sqrt (3)
 
hence another root = 2 + sqrt (3)
 
now
 
sum of roots = b = 4
 
multiplication of roots = c = 4 – 3 = 1
 
hence
 
log b c  = 0
 
b + c = 4 + 1 = 5
 
log cb is not defined